The differential equation



dy/dx= cos(x)(y^2 +6y+8)/(6y+16)




has an implicit general solution of the form F(x,y)=K



In fact, because the differential equation is separable, we can define the solution curve implicitly by a function in the form


F(x,y)=G(x)+H(y)=K


Find such a solution and then give the related functions requested.


F(x,y)=G(x)+H(y)=

1
Expert's answer
2022-03-09T14:33:09-0500

The given equation can be separated as, "\cos x~ dx = \dfrac{6y+16}{(y+2)(y+4)}dy".


Now using partial fractions for the term on the right hand side,


"\begin{aligned}
\frac{6y+16}{(y+2)(y+4)} & = \frac{A}{y+2} + \frac{B}{y+4}\\
6y+16 & =A(y+4) + B(y + 2)\\
\end{aligned}"


Taking "y = -4" gives, "B = 4" and taking "y = -2" gives, "A = 2".

Therefore,


"\cos x~ dx = \left\{\dfrac{2}{y+2} + \dfrac{4}{y+4}\right\}dy"


Integrating both sides, we get

"\begin{aligned}
\sin x &= 2\log (y+2) + 4\log (y+4) \\
\sin x &= \log (y+2)^2 + \log (y+4)^4\\
\sin x + c &= \log \{(y+2)^2 (y+4)^4\}\\
e^{(\sin x) + c} &= (y+2)^2(y+4)^4\\
(y+2)^2(y+4)^4 - e^{\sin x} &= c_1~~~~(\text{where }c_1 = e^c)
\end{aligned}"


Hence, "F(x,y)=G(x)+H(y)= - e^{\sin x} +(y+2)^2(y+4)^4" where


"G(x)= - e^{\sin x}; H(y) = (y+2)^2(y+4)^4".